OpenStudy (anonymous):

WILL FAN AND MEDAL Beth deposited $4000 in a college fund earning 3.5% interest. What was the total interest earned at the end of 12 years? A. $1,166.67 B. $1,680.00 C. $11,666.67 D. $16,800.00

OpenStudy (anonymous):

Assume compounded yearly A = P(1+0.035)^12 is amount after 12 years A = 4000(1.035)^12 = 4000(1.511) = 6044.27 but 4000 of this is Principal, so Interest is I = A-P = 2044.27 Since this does not fit your options, assume simple interest of 0.035 x 12 years I = (0.035)(12)(4000) = 1680 thus B